Using the eliminations method, solve.
4x-2y+8z=6
2x-y+4z=5
6x-3y+12z=7

Answer by ankor@dixie-net.com(12693)

Using the eliminations method, solve.
4x - 2y + 8z = 6
2x - y + 4z = 5
6x - 3y +12z = 7
:
There is no solution here, see what happens when we multiply the last eq by -1:
4x - 2y + 8z = 6
2x - y + 4z = 5
-6x +3y -12z = -7
-------------------- adding eliminates all the varibles